# LLMs syntactically adapt their language use to their conversational partner

### v1

@@ Abstract
Florian Kandra  and Vera Demberg  and Alexander Koller Department of Language Science and Technology Saarland University, Saarbrücken, Germany flka00003@uni-saarland.de {vera, koller}@coli.uni-saarland.de
@@ 1 Introduction
To this end, we create a new dataset of conversations between LLMs in which both LLMs are prompted to initially exhibit different language use. We then measure the dynamics of syntactic language adaptation over the course of the conversations, using a method adapted from the human-human analysis of Reitter and Moore (2014). We find that GPT-4o Hurst et al. (2024) conversations show statistically significant adaptation when comparing syntactic repetitions within conversations against repetitions across conversations, replicating Reitter’s findings for human conversations. We further show this is a continuous process active throughout conversations and conclude by discussing whether these findings demonstrate “human-like” alignment in LLMs.
@@ 2 Background
The only study to date examining adaptation in LLMs Cai et al. (2024) focused on short-term syntactic adaptation. They showed that ChatGPT and Vicuna are more likely to complete a sentence with a double object or a prepositional object when primed with a sentence of the respective type. We extend this research to long conversations with natural sentences rather than carefully constructed one-sentence stimuli.
@@ 3 Measuring human-human adaptation
We follow Reitter in splitting each conversation in a dialogue corpus into two parts. We call the first 49% of each conversation Prime and the last 49% of each conversation Target; the middle 2% are discarded to ensure that we measure long-term adaptation as opposed to short-term priming. On corpora that are not already syntactically annotated, we parse each conversation with the Neural Berkeley Parser [^1] We used the benepar_en3_large model of the benepar python package for parsing and spacy’s en_core_web_md model for tokenization. Kitaev and Klein (2018); Kitaev et al. (2019), to obtain a set of context-free production rules for the Prime and Target section of each conversation, respectively.
Adaptation takes place if rule repetitions are more likely between the Prime and Target of the same conversation (where adaptation is possible), compared to a Prime and Target of different conversations (where no adaptation could have taken place). To make this comparison, we draw two samples for each rule across the Targets of all conversations: one for which we check whether the rule has been uttered by the other speaker in the Prime of the same conversation, and the other for which we check whether the rule has been uttered by a speaker in a random, unrelated conversation. We encode the presence of a rule in a binary variable Prime for each sample, which is $1$ if the rule is present. Another binary variable, SameConv, is used to indicate whether we looked for a prime in the same conversation ($1$) or in a different, random conversation (00). If repetitions are more likely between speakers within conversations, such that we see an effect of SameConv on Prime, we take that as evidence of cross-speaker adaptation.
@@ 3.1 Alignment in human conversations
Table 1: The regression models for the Switchboard corpus (left) and the GPT corpus (right).
We find that SameConv ($\beta=0.228,p<0.001$) has a significant positive effect, replicating Reitter’s findings that humans align syntactically to their partners over the course of a conversation. A thorough analysis of the interactions can be found in Appendix E.
@@ 4 Measuring LLM-LLM adaptation
We follow the same method to analyze syntactic adaptation in conversations of GPT-4o.
@@
We therefore created our own dataset by letting GPT-4o [^2] We used GPT-4o-2024-08-06 with default parameters. converse with itself. [^3] We will make this dataset available upon acceptance. We created 17 different conversational agents with identical system prompts, except for an initial specification of a “language persona” that is unique to each agent. We then generated conversations between pairs of LLM agents by iteratively prompting each of them for the next utterance, including the context of the conversation history. Iterations were stopped, once a conversation surpassed a predefined length threshold. All prompts for managing the conversations and defining the language personas can be found in Appendix A.
To ensure sufficient variety in the agents’ language use, we further generated conversations where each agent conversed with itself. We then calculated how often each syntactic rule was used and normalized these frequencies to create a discrete probability distribution of syntactic rules for each agent. To compare these distributions, we measured their distances using the Jensen-Shannon divergence (JSD). See Figure 5 in Appendix C for details. The results confirm a high degree of syntactic variety, with JSD values of up to 0.69.
@@
We generated 136 conversations by pairing up every conversational agent with every other conversational agent, all on the topic “What makes a day a good day?” Twelve conversations ended in repeating patterns; we excluded them and used the remaining 124 other conversations to form the GPT corpus. The distribution of conversation and utterance lengths closely mirrors that of the Switchboard corpus (cf. Fig. 3 and Fig. 2 in Appendix B).
We ran the analysis described in Section 3.1 on the GPT corpus, taking care not to sample from identical agents in other conversations. Fixed effects, except SameConv, are centered. The model was selected using backward selection. The results are shown in Table 1.
SameConv has a significant positive effect on Prime ($\beta=0.198,\,p<0.001$), showing that there is syntactic adaptation. Interactions are discussed in more detail in Appendix E.
@@
To gain a deeper understanding of the adaptation process performed by the LLM, we performed a fine-grained analysis of adaptation over the course of the conversation. As in the approach above, we directly compared the distributions of syntactic structures used by two different agents; however, this time, we focused on comparing the distributions to see how they evolve throughout a conversation. To obtain reliable estimates of the distributions, we created 520 conversations between agents 5 and 6, a pair of agents with moderate initial JSD (cf. Fig. 5), while keeping the topic the same (cf. Appendix A). Due to repeating patterns, we excluded 14 conversations.
To observe how the similarity of the two agents’ distributions evolves, we split the remaining 506 conversations into sections of 200 words (see Fig. 4 in Appendix B for an overview of the data), and compare the distributions of the two agents for each split. As above, distributions are calculated by normalizing rule frequencies across all 506 conversations. To estimate the variance of these calculations, we perfom them on 100 bootstraps of the data. Each bootstrap consisted of 506 randomly selected conversations, drawn with replacement. We report the means and standard deviations of these 100 JSD values across splits in Fig. 1.
We find that the mutual adaptation of the two LLM agents is a gradual process that persists throughout the conversation. The rate of adaptation is relatively constant, with the strongest adaptation happening in the first split.
@@ 5 Discussion
Throughout the paper, we have avoided using the words “alignment” and “priming” for the LLM’s adaptation process to steer clear of any connotations about human cognitive processes. While we have established that the LLM’s syntax becomes increasingly similar to its conversational partner’s, this does not necessarily mean that this process is driven by a similar underlying mechanism.
An LLM does not maintain an explicit mental model of its interlocutor’s language use and does not make conscious decisions on coordinating it with its interlocutor. Thus it seems inappropriate, under the notion of alignment sketched in Section 2, to explain the LLM’s adaptive behavior as alignment. At the same time, priming effects in humans are usually assumed to impact their language use only in the short term. One conceivable explanation for GPT-4o’s ability to perform long-term adaptation is that it conditions the language it produces on the previous conversation (a mechanism that is similar to priming in humans), but has a much larger capacity than humans for remembering the verbatim conversational context.
This notion goes in support of our second experiment: A gradual adaptation that appears with increased context length underpins the intuition that LLMs can adapt to longer contexts, and that this increases its influence. Different from humans, short-term effects, like those reported in Cai et al. (2024), may therefore be driven by the same principles as long-term adaptation in LLMs. A more detailed analysis would be an interesting avenue of future research.
@@ 6 Conclusion
We showed that GPT-4o can gradually adapt its language use to its conversational partner, to an extent that is similar to what we observe in human-human conversations. This observation goes beyond previous findings, which indicated that an LLM’s language use can be controlled through explicit instructions and influenced by priming from the previous utterance. A more detailed comparison of the mechanisms that humans and LLMs use to achieve such long-term adaptation is an interesting avenue of future work.
@@ 7 Limitations
This work focuses on texts generated with GPT-4o. We decided to use this model, as it is one of the highest performing accessible models. While the findings of Cai et al. (2024) suggest that the results generalize to other models, e.g. Vicuna, further research is needed to confirm this.
Our study concentrates on syntactic structures of the English language. Similar effects may exist for other languages and other linguistic features, also of different modalities (e.g. intonation, speech rate). Furthermore, in this study we controlled for topicality by keeping the topic of all conversations identical. It is unclear whether topicality has an effect on syntactic structures, but there is evidence that lexical choices influence the syntax at least to some extent (lexical boost, Cai et al., 2024). Future work would have to investigate further how the reported effects translate to more diverse conversational settings.
The analysis that we adapt from Reitter and Moore (2014) loses information by encoding the presence of syntactic structures in a binary variable. We hypothesize that this leads to the interaction between ln(Freq):SameConv (see Apeendix E). While the analysis is suitable for capturing adaptation in general, it lacks the sensitivity to account for the occurrence rate of rules in a meaningful way.
@@ 8 Ethical Considerations
We believe that our work is unlikely to have an immediate ethical or societal impact. However, there is potential that the reported effects serve as a footprint of LLM generated texts – we didn’t prompt the model to adapt to the language, but this effect appears inherently. This potentially leads to patterns that are intrinsic to LLMs, which could be leveraged to detect LLM generated texts.
@@ A.2 Language Personas
Figure 3: Statistics of the 136 conversations between agents generated with GPT-4o (GPT Corpus).
Figure 4: Statistics of the 506 conversations between agents 5 and 6 generated with GPT-4o.
@@ Appendix B Dataset Compositions
Statistics of the Switchboard corpus and the conversations generated with GPT-4o are shown in Figures 2 and 3. The composition of the conversations between agents 5 and 6 can be seen in Figure 4. The agents were chosen, as they provide very even turn lengths. This allows for similarly good estimations of the distributions of their used syntactic rules.
The cost for generating all conversations using OpenAI’s API was around 100\$.
Figure 5: Jensen-Shannon divergence values between agents. See section A for an overview of their different language prompts. Agents 14-17 were excluded due to repeating patterns in their conversations.
@@ Appendix C Base Divergence Values between Agents
In our study, we compared the distributions of rules that agents use throughout conversations using the Jensen-Shannon divergence as distance measurement. To place our reported values in context, we provide baseline divergence values between each agent in this appendix. For each agent, we calculated the probability distribution of their uttered rules from 10 conversations with themselves. We counted the number of occurrences of each rule and normalized these frequencies for each agent to create a discrete probability distribution. The topic of all conversations was kept identical: ”What makes a day a good day?” Conversations were created turn by turn and stopped once they surpassed a length of 800 words (see section 4). Agents 14-17 are excluded, as their conversations converged to short repeating patterns. The resulting JSD values between the distributions of each agent are shown in Figure 5.
@@ Appendix D Analysis with all Rules
In our analysis, we exclude rules that have very high frequencies, and those that appear only once. To test whether removing these rules affects overall conclusions, we ran the analysis again using all rules. Results can be found in Table 2 for Switchboard and in Table 3 for the GPT corpus.
@@
Appendix E Interaction Effects
In our analysis of Switchboard we find the interaction between ln(Freq) and SameConv ($\beta=-0.101,p<0.001$), which is similar to the effect size of the same interaction in the GPT corpus ($\beta=-0.146,p<0.001$). The effect could be taken to explain that adaptation appears more for less frequent rules444less than the mean frequency, as the data is centered compared to more frequent rules. In theory, this sounds like a plausible explanation. However, we hypothesize that the effect may primarily come from our binary encoding of Prime: a higher frequent rule reduces the effect that SameConv has on Prime, because it is more likely to appear in other conversations ($SameConv=0$). The binary encoding omits information of the rate at which rules appear, making an increased rate of rules in the Prime and Target invisible. Therefore, the analysis cannot capture adaptation if it manifests as an increase in occurrence rate between the Prime and Target of the same conversation, which reduces sensitivity to more frequent rules.
The effect size of the interaction between the log rule frequency ln(Freq) and the log number of unique rules in the prime ln(Size) differs between GPT ($\beta=0.266,\,p<0.001$) and Switchboard ($\beta=0.068,\,p<0.004$). The larger effect in the GPT corpus suggests a more stochastic mechanism. To illustrate this, we can think about the appearance of syntactic rules, as if they were drawn at random (with replacement) from an underlying probability distribution over rules. Adaptation would mean that drawing a rule increases its probability to be drawn again. In this context, we can think of ln(Size) as the logarithmic number of rules that we draw. For a rule with probability $p$, doubling the Size results in an increase of $p$ by $1-(1-p)^{2}-p=p(1-p)$. It holds that:
$p_{1}(1-p_{1})<p_{2}(1-p_{2}),$
for $0.50<p_{1}<p_{2}<$ 0.5555The most frequent rule has a probability of $0.087$, which is much lower than $0.5$.. Therefore, the probability for high frequency rules increases more with an increase of Size. While the inclusions of ln(Freq) and ln(Size) already account for this effect, the increase in repetition probability also raises the likelihood of further increases due to adaptation, which is captured by the interaction term. The effect therefore highlights the stochasticity in GPT-4o’s behavior.

### v2

@@ Abstract
Florian Kandra, Vera Demberg, Alexander Koller Department of Language Science and Technology Saarland University, Saarbrücken, Germany flka00003@uni-saarland.de {vera, koller}@coli.uni-saarland.de
@@ 1 Introduction
To this end, we create a new dataset of conversations between LLMs in which both LLMs are prompted to initially exhibit different language use. We then measure the dynamics of syntactic language adaptation over the course of the conversations, using a method adapted from the human-human analysis of Reitter and Moore (2014). We find that GPT-4o Hurst et al. (2024) and Llama-3-8B Grattafiori et al. (2024) conversations show statistically significant adaptation when comparing syntactic repetitions within conversations against repetitions across conversations, replicating Reitter’s findings for human conversations. We further show this is a continuous process active throughout conversations and conclude by discussing whether these findings demonstrate “human-like” alignment in LLMs.
Our code and data are publicly available [^1] https://github.com/coli-saar/llm-language-accomodation.
@@ 2 Background
Cai et al. (2024) examined syntactic adaptation in LLMs, with a focus on short-term priming. They showed that ChatGPT and Vicuna are more likely to complete a sentence with a double object or a prepositional object when primed with a sentence of the respective type. We extend this research to long conversations with natural sentences rather than carefully constructed one-sentence stimuli.
@@ 3 Measuring human-human adaptation
[…]
Figure 1: Phrase Structure Tree and Extracted Production Rules
We follow Reitter in splitting each conversation in a dialogue corpus into two parts. We call the first 49% of each conversation Prime and the last 49% of each conversation Target; the middle 2% are discarded to ensure that we measure long-term adaptation as opposed to short-term priming. On corpora that are not already syntactically annotated, we parse each conversation with the Neural Berkeley Parser [^2] We used the benepar_en3_large model of the benepar python package for parsing and spacy’s en_core_web_md model for tokenization. Kitaev and Klein (2018); Kitaev et al. (2019), to obtain a set of context-free production rules for the Prime and Target section of each conversation, respectively (see Fig. 1).
Figure 2: Sampling process to analyse syntactic alignment. Samples are drawn by checking rule occurences in the same conversation and in different random conversations.
Adaptation takes place if rule repetitions are more likely between the Prime and Target of the same conversation (where adaptation is possible), compared to a Prime and Target of different conversations (where no adaptation could have taken place). In order to fit a model that can find such an effect, we sample data from our conversations in the following way: For each rule across the Targets of all conversations, we draw two samples, one for which we check whether that rule has been uttered within the same conversation, but by the other speaker, and one sample for which we check whether that rule has been uttered by a random speaker of a random other conversation. Figure 2 depicts this process. For every sample, we encode whether a rule was found this way in a binary variable Prime. Another binary variable, SameConv, is used to indicate whether we looked for a prime in the same conversation ($1$) or in a different, random conversation (00). If repetitions are more likely between speakers within conversations, such that we see an effect of SameConv on Prime, we take that as evidence of cross-speaker adaptation.
@@ 3.1 Alignment in human conversations
Llama Corpus
$\beta$ SE $z$ $p>|z|$
-2.016 0.049 -41.0 0.000
1.333 0.030 44.5 0.000
0.505 0.052 9.6 0.000
0.825 0.147 5.6 0.000
-0.252 0.032 -7.8 0.000
-0.005 0.088 -0.06 0.952
Table 1: The regression models for the Switchboard corpus (left) and the GPT corpus (middle) and LLama corpus (right). Effects show high significance except for the interaction between ln(Freq) and ln(Size) in the Llama corpus.
We find that SameConv ($\beta=0.228,p<0.001$) has a significant positive effect, replicating Reitter’s findings that humans align syntactically to their partners over the course of a conversation.
@@ 4 Measuring LLM-LLM adaptation
We follow the same method to analyze syntactic adaptation in conversations of GPT-4o and Llama-3-8B.
@@
We therefore created our own dataset by letting GPT-4o and Llama-3-8B [^3] We used GPT-4o-2024-08-06 and Meta-Llama-3-8B-Instruct with default parameters. converse with themselves. We created 17 different conversational agents with identical system prompts, except for an initial specification of a “language persona” that is unique to each agent. We then generated conversations between pairs of LLM agents by iteratively prompting each of them for the next utterance, including the context of the conversation history. Iterations were stopped, once a conversation surpassed a predefined length threshold. All prompts for managing the conversations and defining the language personas can be found in Appendix A.
To ensure sufficient variety in the agents’ language use, we further generated conversations where each agent conversed with itself. We then calculated how often each syntactic rule was used and normalized these frequencies to create a discrete probability distribution of syntactic rules for each agent. To compare these distributions, we measured their distances using the Jensen-Shannon divergence (JSD). See Figures 9 and 10 in Appendix C for details. The results confirm a high degree of syntactic variety, with JSD values of up to 0.69 for GPT-4o and 0.70 for Llama-3-8B.
@@
We generated 136 conversations for each model by pairing up every conversational agent with every other conversational agent, all on the topic “What makes a day a good day?” In GPT conversations, twelve of them ended in repeating patterns (see Appendix E) for Llama only one; we excluded them and used the remaining 124 and 135 conversations to form the GPT and Llama corpora respectively. The distributions of conversation and utterance lengths closely mirror that of the Switchboard corpus (cf. Fig. 5, Fig. 6 and Fig. 4 in Appendix B).
We ran the analysis described in Section 3.1 on the corpora. Because agents appear in multiple different conversations, we took care not to sample from identical agents from other conversations. Fixed effects, except SameConv, are centered. The models were selected using backward selection. The results are shown in Table 1.
For both Llama and GPT, SameConv has a significant positive effect on Prime ($\beta=0.198,\,p<0.001$ and $\beta=0.364,\,p<0.001$), showing that there is syntactic adaptation.
@@
To gain a deeper understanding of the adaptation process performed by the LLMs, we performed a fine-grained analysis of adaptation over the course of the conversation. To do so, we again directly compared the distributions of syntactic structures used by two different agents; however, this time, we focused on comparing the distributions to see how they evolve throughout a conversation. To obtain reliable estimates of the distributions, we created 520 conversations between agents 5 and 6, a pair of agents with moderate initial JSD (cf. Fig. 9, 10), while keeping the topic the same (cf. Appendix A). Due to repeating patterns, we excluded 14 conversations of GPT-4o and 7 of the Llama model.
To observe how the similarity of the two agents’ distributions evolves, we split the remaining conversations into sections of 200 words (see Fig. 7, 8 in Appendix B for an overview of the data), and compare the distributions of used rules by the two agents for each split. Again, we obtain these probability distributions by normalizing rule frequencies for each split across conversations. To estimate the variance of these calculations, we perform them on 100 bootstraps of the data. Each bootstrap is a collection of conversations that is drawn with replacement from the original GPT and Llama conversations, therefore resulting in 100 collections of 506 and 513 randomly drawn conversations for each model. We report the means and standard deviations of these 100 JSD values across splits in Fig. 3.
We find that the mutual adaptation of the two LLM agents is a gradual process that persists throughout the conversation for both Llama-3-8B and GPT-4o. The rate of adaptation is relatively constant, with the strongest adaptation happening in the first split.
@@ 5 Discussion
Throughout the paper, we have avoided using the words “alignment” and “priming” for the LLMs’ adaptation process to steer clear of any connotations about human cognitive processes. While we have established that the LLMs’ syntax becomes increasingly similar to their conversational partner’s, this does not necessarily mean that this process is driven by a similar underlying mechanism.
An LLM does not maintain an explicit mental model of its interlocutor’s language use and does not make conscious decisions on coordinating it with its interlocutor. Thus it seems inappropriate, under the notion of alignment sketched in Section 2, to explain the LLMs’ adaptive behavior as alignment. At the same time, priming effects in humans are usually assumed to impact their language use only in the short term. One conceivable explanation for GPT-4o’s and Llama-3-8B’s ability to perform long-term adaptation is that they condition the language they produce on the previous conversation (a mechanism that is similar to priming in humans), but have a much larger capacity than humans for remembering the verbatim conversational context.
The findings of our second experiment also support this idea: A gradual adaptation that appears with increased context length underpins the intuition that LLMs can adapt to longer contexts, and that longer context correlates with an increased influence on the production of syntactic structures. Different from humans, short-term effects, like those reported in Cai et al. (2024), may therefore be driven by the same principles as long-term adaptation in LLMs. A more detailed analysis would be an interesting avenue of future research.
@@ 6 Conclusion
We showed that GPT-4o and Llama-3-8B can gradually adapt their language use to their conversational partner, to an extent that is similar to what we observe in human-human conversations. This observation goes beyond previous findings, which indicated that an LLM’s language use can be controlled through explicit instructions and influenced by priming from the previous utterance. A more detailed comparison of the mechanisms that humans and LLMs use to achieve such long-term adaptation is an interesting avenue of future work.
7 Acknowledgements
This research was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), which supported the work of FK – Project KO 2916/3-1 – and VD – project number 389792660 – TRR 248 – CPEC, see https://perspicuous-computing.science.
Florian Kandra would like to thank Albert Kandra, who helped with extensive analytical conversations during the preparation of this paper.
@@ 8 Limitations
This work focuses on texts generated with GPT-4o and Llama-3-8B. We decided to use these models, as they cover a range between one of the highest performing accessible models and an open sourced, but smaller and less performant model. The findings suggest that syntactic adaptation generalizes to other models as well, which is supported by the findings of Cai et al. (2024). Further research could explore additional models to investigate the underlying factors that may influence syntactic adaptation.
Our study concentrates on syntactic structures of the English language. LLMs may exhibit similar behavior for other languages and other linguistic features, also of different modalities (e.g. intonation, speech rate). It would be interesting to investigate this in future work.
Furthermore, in this study we controlled for topicality by keeping the topic of all conversations identical. It is unclear whether topicality has an effect on syntactic structures, but there is evidence that lexical choices influence the syntax at least to some extent (lexical boost, Cai et al., 2024). To what extent topicality can have an effect on syntactic choices in LLMs is left as an avenue for future research.
Moreover, our study currently only focuses on LLM-LLM conversations. It would be interesting to see how these effects impact human-LLM conversations, especially given the societal impact of human-LLM interaction.
The analysis that we adapt from Reitter and Moore (2014) loses information by encoding the presence of syntactic structures in a binary variable. While the analysis is suitable for capturing adaptation in general, it lacks the sensitivity to account for the occurrence rate of rules in a meaningful way.
@@ 9 Ethical Considerations
We believe that our work is unlikely to have an immediate ethical or societal impact. However, there is potential that the reported effects serve as a footprint of LLM generated conversations – we didn’t prompt the model to adapt to the language, but this effect appears inherently. This potentially leads to patterns that are intrinsic to LLMs, which could be leveraged to detect LLM generated conversations.
@@ A.2 Language Personas
Figure 5: Statistics of the 124 conversations between agents generated with GPT-4o (GPT Corpus).
Figure 6: Statistics of the 135 conversations between agents generated with Llama-3-8B (Llama Corpus).
Figure 7: Statistics of the 506 conversations between agents 5 and 6 generated with GPT-4o. Utterance lengths (right) are averaged across all conversations for each split and for both agents.
Figure 8: Statistics of the 513 conversations between agents 5 and 6 generated with Llama-3-8B. Utterance lengths (right) are averaged across all conversations for each split and for both agents.
@@ Appendix B Dataset Compositions
Statistics of the Switchboard corpus and the conversations generated with GPT-4o and Llama-3-8B are shown in Figure 4, Figure 5, and Figure 6 respectively. The compositions of the conversations between agents 5 and 6 for GPT-4o and Llama-3-8B can be seen in Figures 7 and 8. The agents were chosen, as they provide very even utterance lengths across splits. This allows for similarly good estimations on their rule probability distributions.
The cost for generating all conversations using OpenAI’s API was around 100\$. Generating the Llama conversations took $\approx$ 7 GPU hours on a single NVIDIA H100.
Figure 9: Jensen-Shannon divergence values between agents. Distributions of syntactic rules are taken from 10 GPT-4o conversations of all agents with themselves (see Appendix C). Appendix A gives an overview of their different language prompts. Agents 14-17 were excluded due to repeating patterns in their conversations (see Appendix E).
Figure 10: Jensen-Shannon divergence values between agents. Distributions of syntactic rules are taken from 10 Llama-3-8B conversations of all agents with themselves (see Appendix C). Appendix A gives an overview of their different language prompts.
@@ Appendix C Base Divergence Values between Agents
In our study, we compared the distributions of rules that agents use throughout conversations using the Jensen-Shannon divergence as distance measurement. To place our reported values in context, we provide baseline divergence values between each agent on an additional set of conversations that we present here. For each agent, we calculated their probability distribution of uttered rules from 10 conversations with themselves. The topic of all conversations was again kept identical: ”What makes a day a good day?” Conversations were created turn by turn and stopped once they surpassed a length of 800 words (see Section 4). Probability distributions are taken to be the normalized frequencies of rule occurrences in the 10 conversations of each agent. Conversations were created using GPT-4o and Llama-3-8B. In the GPT conversations, agents 14-17 ended in repeating patterns and were excluded as a result (see Appendix E). The resulting JSD values of pairwise comparisons between each agent are shown in Figure 9 for GPT-4o and Figure 10 for Llama.
@@ Appendix D Analysis with all Rules
In our analysis, we exclude rules that have very high frequencies, and those that appear only once. To test whether removing these rules affects overall conclusions, we ran the analysis again on the GPT corpus using all rules. Results can be found in Table 2 for Switchboard and in Table 3 for the GPT corpus.
@@
Appendix E Repeating Patterns in Conversations
Some conversations between Llama-3-8B or GPT-4o ended in repeating patterns. These were short, repeated statements from both agents, including repeated expressions of gratitude, farewells, or affirmational remarks. Examples are provided in Figure 11.
{adjustwidth}
3pt3pt
GPT Agent3 with itself
Agent3: Indeed. Action is the essential link in that chain.
Agent3: Correct. It’s crucial for progress.
Agent3: Absolutely. Without it, progress stalls.
Agent3: Yes. Action keeps momentum alive.
Agent3: Precisely. It fuels continuous movement forward.
Agent3: Correct. It sustains growth and development.
GPT Agent3 and Agent5
Agent3: Later!
Agent5: Till next time!
Agent3: Till then!
Agent5: See you soon!
Agent3: See you!
Figure 11: Repeating patterns in GPT conversations. An affirmative pattern is shown on the top, a loop of parting expressions on the bottom.